# Quasiconservation Laws and Suppressed Transport in Weakly Interacting Localized Models

### v2

@@ II.2 Local Integrals of Motion
We note that $\tilde{n}_{\alpha}^{(0)}\equiv\hat{d}^{\dagger}_{\alpha}\hat{d}_{\alpha}$, which represent the occupations of the localized orbitals, are conserved quantities of $\hat{H}_{0}$, so that $[H_{0},\tilde{n}_{\alpha}^{(0)}]=0$. This gives $\hat{H}_{0}$ a noninteracting integrability, where $\hat{H}_{0}$ can be diagonalized with these occupations.
@@ III Quasi-LIOMs
$[\hat{H}_{0},\hat{d}^{\dagger}_{\alpha}\hat{d}_{\alpha}]=[\hat{H}_{0},\tilde{n}^{(0)}_{\alpha}]=0.$ (8)
Here, $\alpha=1,\dots,L$, and for any $\alpha,\beta$, $[\tilde{n}^{(0)}_{\alpha},\tilde{n}^{(0)}_{\beta}]=0$. These conservation laws are broken when the perturbation given in Eq. (3) is added. We consider the problem of finding corrections $\tilde{n}^{(1)}_{\alpha}$ to $\tilde{n}^{(0)}_{\alpha}$, such that the quantity $\tilde{n}^{(0)}_{\alpha}+\lambda\tilde{n}^{(1)}$ still commutes with $H_{0}+\lambda V$ up to second order in $\lambda$:
@@
In this work, we call the quantity $\tilde{n}^{(0)}_{\alpha}+\lambda\tilde{n}^{(1)}$ a quasi-LIOM, following the naming convention of quasi-charges. The quasi-conservation law above is equivalent to the relation:
@@
[…]
One formal solution to the problem of finding quasi-LIOMs is given by the Adiabatic Gauge Potential (AGP), $\hat{X}$. We split the perturbation $\hat{V}$ into a diagonal and off-diagonal part,
@@
Note that this operator is defined up to additive terms that commute with $H_{0}$.
Using $\hat{X}$, we can define the corrections $\tilde{n}^{(1)}_{\alpha}$:
$\tilde{n}^{(1)}_{\alpha}=i\left[\hat{X},\tilde{n}^{(0)}_{\alpha}\right]\leavevmode\nobreak\ .$ (13)
A simple application of the Jacobi identity $[[\hat{A},\hat{B}],\hat{C}]+[[\hat{B},\hat{C}],\hat{A}]+[[\hat{C},\hat{A}],\hat{B}]=0$ with $\hat{A},\hat{B},\hat{C}$ being $\hat{X},\tilde{n}^{(0)},\hat{H}_{0}$, respectively, shows that the form of LIOM corrections in Eq. (13) satisfies Eq. (10).
@@ III.2 Methods
In this work, we study most quantities (except for noninteracting charge transport) by computing their Frobenius norms, which are computationally more tractable than the operator norms. Formulas for Frobenius norms of four-fermionic operators can be found in Appendix B. For exponentially localized operators, both Frobenius and operator norms are expected to be finite in the thermodynamic limit. We are mostly interested in assessing whether $\tilde{n}_{\alpha}^{(1)}$ is localized. Therefore, by verifying whether its Frobenius norm is finite and bounded, we obtain strong evidence of its localization.
@@
For the disordered model, calculations are presented with a distribution over independent and randomly generated disorder realizations. We also obtain distributions for the quasiperiodic model. Although the model is deterministic, our calculations found that changing the phase shift $\delta$ affects the results. To avoid commensuration issues, we found it useful to sample over these variations with a distribution over the phase $\delta$, spanning 500 evenly spaced phases between $0<\delta<1/k$, which covers one period of the potential. This setup also allows subsequent analysis of interaction corrections to charge transport in both models.
@@ III.3.1 Disordered Model AGP
Figure 1: Analysis of the AGP Frobenius norm in the interacting Anderson Model. All three figures were generated based on 500 disorder realizations and open boundary conditions. (a) Calculated probability distributions of $\ln\|\hat{X}\|_{F}$ with disorder strength $W=2.1$. Each colored curve represents varying system sizes $L$, from $L=20$ to $L=100$ in steps of 101010. (b) Same data as in (a) but with the vertical scale shown on a log scale. The red line represents a polynomial $\|\hat{X}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\hat{X}\|_{F}$—see text around Eq. (21) for details on the form plotted. (c) Median of the AGP norm as a function of system size $L$. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in step sizes of 0.5. Here, dashed lines represent power-law fits to the data, $\|\hat{X}\|_{F}=AL^{D}+C$ with power-law exponents $D$. In order of increasing disorder strength, we find $D=2.33,2.33,1.99,1.71,1.45,1.38,1.34$, consistent with expected eventual linear in $L$ scaling for a strictly localized system (see text in Sec. III.3.1 for discussion). The error bars on the median curves, estimated using a bootstrapping method, are very small and are therefore omitted in this and subsequent plots for clarity.
@@
The most significant feature of the calculated probability distribution $P\left(\|\hat{X}\|_{F}\right)$ are the long inverse square power-law tails $\sim\frac{1}{\|\hat{X}\|_{F}^{2}}$. The curve corresponding to $P\left(\|\hat{X}\|_{F}\right)\sim\|\hat{X}\|_{F}^{-2}$ is shown as a comparison in Fig. 1(b). We note here that a power-law $\sim\!x^{-p}$ tail of the probability distribution $P(x)$, with $p\in\mathbb{R}$, translates to the tail
@@
The inverse square tail arises naturally from the distribution of the denominators $x\equiv\frac{1}{\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\delta}-\epsilon_{\gamma}}$. If one approximates the eigenvalues $\epsilon_{\alpha}$ as being randomly sampled from a uniform distribution between $-W^{\prime}$ and $W^{\prime}$ (the eigenvalues do not exactly follow a uniform distribution, but it is a reasonable approximation), then the distribution of the reciprocal variable $x\equiv 1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\delta}-\epsilon_{\gamma})$ follows $P_{1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\delta}-\epsilon_{\gamma})}(x)\approx\frac{1}{x^{2}}$ at very large $x$. Details of the derivation can be found in Appendix D. Even though many of the numerators $V_{\alpha\beta\gamma\delta}$ are very small due to the Anderson localization of the eigenstates $\phi_{\alpha}(j)$, we still expect $O(L)$ of them to be $O(1)$. Therefore, the distribution of the Frobenius norm of $\hat{X}$ will also present a tail $\propto 1/\|\hat{X}\|_{F}^{2}$ at large $\|\hat{X}\|_{F}$, as seen in the data. As a result, we expect these tails to exist for any disorder strength $W$. Indeed, we performed similar calculations for disorder strengths much higher than the ones presented in this paper ($W=6$) and verified that these inverse power-law tails exist even at high disorders. This remark applies to other quantities calculated in this work that regard the disordered Anderson model.
@@ III.3.2 Quasiperiodic Model AGP
When the states are delocalized at low disorders for $h<2$, one also observes polynomial growth of $\|\hat{X}\|_{F}$. Curiously, in this delocalized regime, $\mathrm{med}\|\hat{X}\|_{F}$ grows non-monotonically, and $\mathrm{med}\|\hat{X}\|_{F}$ spikes at certain system sizes. In the case with $k=\sqrt{2}$, this occurs at $L=40$. These spikes shift to different system sizes when the wavenumber is changed to different irrational numbers, which we have separately checked for $k=\phi$, $k=\sqrt{3}$, and $k=\sqrt{6}$. Here, $\phi$ is the golden ratio. Anomalous behavior of this kind is known to arise in the quasiperiodic model at certain integer distances. For example, Ref. [37] observed that single-particle resonances are enhanced at orbitals separated by distances equal to the Fibonacci numbers $F_{n}$, when one picks $k=\phi$. Similarly, Ref. [74] reported suppressed interaction between LIOMs separated at the Fibonacci numbers for $k=\phi$, and the corresponding integer sequences that relate to the silver and bronze ratios. In our case, we find a more general condition that explains these anomalous system sizes, encompassing previous observations: they occur when the underlying potential is nearly commensurate with the lattice. For open boundary conditions, this corresponds to when $k(L+1)\approx n$ for some $n\in\mathbb{Z}$, so $k\approx\frac{n}{L+1}$ for some integer $n$. For example, in the calculation presented in Fig. 2(c), which was done for $k=\sqrt{2}$, there is a notable spike at $L=40$, and $\sqrt{2}\times 41\approx 58$. This observation helps explain why corresponding integer sequences arise for a given irrational number $k$, an example being the Fibonacci numbers when one chooses $k=\phi$, the golden ratio. Here, we note that 414141 is a member of the integer sequence $\frac{1}{2}Q_{n}$, where $Q_{n}$ are the Pell-Lucas numbers. Just as successive ratios of the Fibonacci numbers form the best rational approximations for $\phi$ ($\lim_{n\rightarrow\infty}F_{n+1}/F_{n}=\phi$), similarly, $\lim_{n\rightarrow\infty}\frac{1}{2}Q_{n}/P_{n}=\sqrt{2}$, where $P_{n}$ here refers to the Pell numbers. Thus, if $L+1=\frac{1}{2}Q_{n}$ for some $n$, then $(L+1)\times\sqrt{2}\approx m\in\mathbb{Z}$, leading to the potential being nearly commensurate with the lattice. We confirm that for $k=\phi$, these spikes occur at system sizes corresponding to the Fibonacci numbers and other integers $m$ that happen to satisfy $m\times k\approx n$. This underlying commensurability is likely the culprit that leads to increased resonances and other strange behavior in the spectrum, leading to a spike in the quantities calculated here. This would be expected as a periodic system exhibits vastly different physics from an incommensurate one.
@@ III.4.1 LIOM Corrections: Anderson Model
Figure 3: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $|\tilde{n}_{\alpha}^{(1)}|_{F}$, in the weakly interacting Anderson model. Each plot displays all orbitals $\alpha=1,\dots,L$ across the 500 disorder realizations used to generate the data shown in this figure. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ at $W=0.6,1.6,2.6,3.1$ and for $\alpha=1,\dots,L$. Within each panel, each colored curve represents a different system size. We show calculations between $L=20$ and $L=100$ in steps of 101010. The inset in (d) shows the same data as in the main (d) panel, but with the vertical scale shown on log-scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (21) for details on the form plotted. The convergence of the distributions for large enough $W$ and $L$ indicates the presence of an extensive number of quasiconservation laws after interactions are added, even in the thermodynamic limit. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in steps of 0.50.50.5. One observes that the curves appear to saturate for large $L$, indicating that at least half of the LIOMs are conserved to second order when interactions are added. See Sec. III.4.1 for a full discussion of the results.
@@
We present numerical calculations of the Frobenius norm of the LIOM corrections, $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$, in Fig. 3, for the weakly–interacting Anderson model. For each $L$, we sample 500500500 disorder realizations and record $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ for every orbital $\alpha=1,2,...,L$ in each disorder realization, giving a total of $500\times L$ datapoints in each distribution.
@@
The data in Fig. 3(a)–(d) suggest that the LIOM corrections converge to a fixed distribution in the thermodynamic limit. At low disorder, finite-size effects play a role, and the eigenstates appear extended. In these cases, the distributions initially shift rightward, but appear to converge at large $L$. At high disorder, the wavefunctions are localized and the distributions are converged, even at low $L$. Additionally, Fig. 3(e) shows how the median of these distributions grows with $L$. Here, we find that at low disorder, the median of $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ grows as a polynomial with $L$. At higher disorders, we see similar growth of $\mathrm{med}\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ for low $L$, but the curves appear to saturate with increasing system size, suggesting that at least half of the noninteracting LIOMs admit bounded corrections in the thermodynamic limit for high enough disorder.
@@
Furthermore, in Fig. 4, we note that the $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ vs $L$ curves collapse when the x-axis is appropriately scaled by the participation ratio of the Anderson model, which scales approximately as $\mathrm{PR}(W)\propto W^{-1.78}$ in the range of $W$ considered here (see Appendix E for a more detailed explanation) and when the y-axis is scaled phenomenologically with $\sim\mathrm{PR}^{1.7}$. Here, the participation ratio is used as a proxy [^3] Some caution is needed in identifying the median participation ratio with the localization length in this analysis. First of all, both the participation ratio and the localization length depend on the energy of the noninteracting orbital, while our analysis does not consider the energy resolution of the LIOM corrections. Moreover, the power-law dependence of the localization length on the disorder strength is expected not to hold at large $W$, where a more refined ansatz is needed [75]. See Appendix E for another discussion. for the localization length $\xi$. The collapse of these curves suggests that at low disorders, the growth of $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is a finite-size effect. If the system size $L$ exceeds the localization length, for any disorder strength $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is expected to saturate to a finite value. Put together, these results suggest that even in the thermodynamic limit, a fraction of noninteracting LIOMs have finite and small corrections in the presence of interactions.

### v3

@@ II.2 Local Integrals of Motion
We note that $\tilde{n}_{\alpha}^{(0)}\equiv\hat{d}^{\dagger}_{\alpha}\hat{d}_{\alpha}$, which represent the occupations of the localized orbitals, are conserved quantities of $\hat{H}_{0}$, so that $H_{0},\tilde{n}_{\alpha}^{(0)}]=0$. This gives $\hat{H}_{0}$ a noninteracting integrability, where $\hat{H}_{0}$ can be diagonalized with these occupations.
@@ III Approximate LIOMs
$\hat{H}_{0},\hat{d}^{\dagger}_{\alpha}\hat{d}_{\alpha}]=[\hat{H}_{0},\tilde{n}^{(0)}_{\alpha}]=0.$ (8)
Here, $\alpha=1,\dots,L$, and for any $\alpha,\beta$, $\tilde{n}^{(0)}_{\alpha},\tilde{n}^{(0)}_{\beta}]=0$. These conservation laws are broken when the perturbation given in Eq. (3) is added. We consider the problem of finding corrections $\tilde{n}^{(1)}_{\alpha}$ to $\tilde{n}^{(0)}_{\alpha}$, such that the quantity $\tilde{n}^{(0)}_{\alpha}+\lambda\tilde{n}^{(1)}$ still commutes with $H_{0}+\lambda V$ up to second order in $\lambda$:
@@
The quantity $\tilde{n}^{(0)}_{\alpha}+\lambda\tilde{n}^{(1)}$ is an “approximate LIOM”. The quasi-conservation law above is equivalent to the relation:
@@
One important note is that one could, in principle, analyze higher $l$-order LIOM corrections $\tilde{n}^{(l)}_{\alpha}$, such that the quasiconservation law holds up to $O(\lambda^{l})$ for some $l\in\mathbb{Z^{+}}$. These corrections will be (2l+2)(2l+2)–fermion terms, and their numerical calculations will scale at least as $L^{2l+2}$, becoming very expensive. Furthermore, the energy resonances that arise from perturbation theory will grow increasingly complex. As a result, we do not attempt higher-order calculations in this work.
One formal solution to the problem of finding approximate LIOMs is given by the Adiabatic Gauge Potential (AGP), $\hat{X}$. We split the perturbation $\hat{V}$ into a diagonal and off-diagonal part,
@@
Note that this operator is defined up to additive terms that commute with $H_{0}$.  [^3] Conventionally, the AGP is defined as the generator of adiabatic transformations of the eigenstates $|n(\lambda)\rangle$ of a Hamiltonian $\hat{H}(\lambda)$ [71, 58, 72]. In our notation (note that the convention used in some of the literature has the opposite sign for $\hat{X}$): $\hat{X}|n(\lambda)\rangle=-i\partial_{\lambda}|n(\lambda)\rangle.$ (13) The relation between our definition of the AGP and Eq. (13) is seen by observing that Eq. (13) obeys the relation $\partial_{\lambda}\hat{H}(\lambda)=i[\hat{X},\hat{H}(\lambda)]+\sum_{n}\partial_{\lambda}E(\lambda)|n(\lambda)\rangle\langle n(\lambda)|.$ (14) One can verify this property by assuming the definition Eq. (14), then evaluating the left-hand side $\partial_{\lambda}\hat{H}=\partial_{\lambda}\sum_{n}E_{n}|n\rangle\langle n|$. If one sets $\lambda=0$ on the right-hand side of Eq. (14), then one recovers Eq. (12). Specifically, in our case, $\partial_{\lambda}\hat{H}(\lambda)=\hat{V}$, the first term on the right-hand side is $i[\hat{X},\hat{H}_{0}]=\hat{V}_{\mathrm{off-diag}}$ and the second term on the right-hand side is $\hat{V}_{\mathrm{diag}}$.
Now, using $\hat{X}$, we can define the corrections $\tilde{n}^{(1)}_{\alpha}$:
[…]
A simple application of the Jacobi identity $\hat{A},\hat{B}],\hat{C}]+[[\hat{B},\hat{C}],\hat{A}]+[[\hat{C},\hat{A}],\hat{B}]=0$ with $\hat{A},\hat{B},\hat{C}$ being $\hat{X},\tilde{n}^{(0)},\hat{H}_{0}$, respectively, shows that the form of LIOM corrections in Eq. (15) satisfies Eq. (10).
@@ III.2 Methods
In this work, we study most quantities (except for noninteracting charge transport) by computing their Frobenius norms, which are computationally more tractable than the operator norms. Here, we define the Frobenius norm of an operator $\hat{O}$ as the normalized Hilbert-Schmidt inner-product
$\|\hat{O}\|_{F}^{2}=\frac{1}{\mathcal{D}}\mathrm{Tr}[\hat{O}^{\dagger}\hat{O}$], (23)
where $\mathcal{D}$ is the total Hilbert space dimension. We are interested in the normalized Frobenius norm as this results in localized operators having $O(1)$ norm. Formulas for Frobenius norms of four-fermionic operators can be found in Appendix B. For exponentially localized operators, both Frobenius and operator norms are expected to be finite in the thermodynamic limit. Here, we are mostly interested in assessing whether $\tilde{n}_{\alpha}^{(1)}$ is localized. Therefore, by verifying whether its Frobenius norm is finite and bounded, we obtain strong evidence of its localization.
@@
For the disordered model, calculations are presented with a distribution over independent and randomly generated disorder realizations. We also obtain distributions for the quasiperiodic model. Although the model is deterministic, our calculations found that changing the phase shift $\delta$ affects the results. To avoid commensuration issues, we found it useful to sample over these variations with a distribution over the phase $\delta$, spanning 500 evenly spaced phases between $0\leq\delta<1/k$, which covers one period of the potential. This setup also allows subsequent analysis of interaction corrections to charge transport in both models.
@@ III.3.1 Disordered Model AGP
Figure 1: Analysis of the AGP Frobenius norm in the interacting Anderson Model. All three figures were generated based on 500 disorder realizations and open boundary conditions. (a) Calculated probability distributions of $\ln\|\hat{X}\|_{F}$ with disorder strength $W=2.1$. Each colored curve represents varying system sizes $L$, from $L=20$ to $L=100$ in steps of $10$. (b) Same data as in (a) but with the vertical scale shown on a log scale. The red line represents a polynomial $\|\hat{X}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\hat{X}\|_{F}$—see text around Eq. (24) for details on the form plotted. (c) Median of the AGP norm as a function of system size $L$. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in step sizes of 0.5. Here, dashed lines represent power-law fits to the data, $\|\hat{X}\|_{F}=AL^{D}+C$ with power-law exponents $D$. In order of increasing disorder strength, we find $D=2.33,2.33,1.99,1.71,1.45,1.38,1.34$, consistent with expected eventual linear in $L$ scaling for a strictly localized system (see text in Sec. III.3.1 for discussion). The error bars on the median curves, estimated using a bootstrapping method, are very small and are therefore omitted in this and subsequent plots for clarity.
@@
The most significant feature of the calculated probability distribution of the norm of the AGP, $P\left(\|\hat{X}\|_{F}\right)$, are the long inverse square power-law tails $\sim\frac{1}{\|\hat{X}\|_{F}^{2}}$. The curve corresponding to $P\left(\|\hat{X}\|_{F}\right)\sim\|\hat{X}\|_{F}^{-2}$ is shown as a comparison in Fig. 1(b). We emphasize that this is the tail for the probability distribution of $\|\hat{X}\|_{F}$, not for the plotted distribution of $\ln(\|\hat{X}\|_{F})$. The conversion between the two is straightforward: a power-law $\sim\!x^{-p}$ tail of the probability distribution $P(x)$, with $p\in\mathbb{R}$, translates to the tail
@@
The inverse square tail arises naturally from the distribution of the denominators $x\equiv\frac{1}{\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\delta}-\epsilon_{\gamma}}$. If one approximates the eigenvalues $\epsilon_{\alpha}$ as being randomly sampled from a uniform distribution between $W^{\prime}$ and $W^{\prime}$ (the eigenvalues do not exactly follow a uniform distribution, but it is a reasonable approximation), then the distribution of the reciprocal variable $x\equiv 1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\delta}-\epsilon_{\gamma})$ follows $P_{1/(\epsilon_{\alpha}+\epsilon_{\beta}-\epsilon_{\delta}-\epsilon_{\gamma})}(x)\approx\frac{1}{x^{2}}$ at very large $x$. Details of the derivation can be found in Appendix D. Even though many of the numerators $V_{\alpha\beta\gamma\delta}$ are very small due to the Anderson localization of the eigenstates $\phi_{\alpha}(j)$, we still expect $O(L)$ of them to be $O(1)$. Therefore, the distribution of the Frobenius norm of $\hat{X}$ will also present a tail $\propto 1/\|\hat{X}\|_{F}^{2}$ at large $\|\hat{X}\|_{F}$, as seen in the data. As a result, we expect these tails to exist for any disorder strength $W$. Indeed, we performed similar calculations for disorder strengths much higher than the ones presented in this paper ($W=6$) and verified that these inverse power-law tails exist even at high disorders. This remark applies to other quantities calculated in this work that regard the disordered Anderson model.
@@
Finally, we briefly comment on the dependence of the distributions of the AGP norms on $L$. Consistent with the power-law scaling of the median of the AGP norms observed in panel (c) of Fig. 1, the probability distributions in panel (b) remain unconverged even at the largest system sizes studied, and shift rightwards at a linear pace with increasing $L$. For clarity, we emphasize that the plotted quantity is actually the distribution of $\ln\|\hat{X}\|_{F}$, chosen for visibility. This transformation can create the misleading visual impression that the distributions are converging, whereas in reality they are not.
@@ III.3.2 Quasiperiodic Model AGP
When the states are delocalized at low disorders for $h<2$, one also observes polynomial growth of $\|\hat{X}\|_{F}$. Curiously, in this delocalized regime, $\mathrm{med}\|\hat{X}\|_{F}$ grows non-monotonically, and $\mathrm{med}\|\hat{X}\|_{F}$ spikes at certain system sizes. In the case with $k=\sqrt{2}$, this occurs at $L=40$. These spikes shift to different system sizes when the wavenumber is changed to different irrational numbers, which we have separately checked for $k=\phi$, $k=\sqrt{3}$, and $k=\sqrt{6}$. Here, $\phi$ is the golden ratio. Anomalous behavior of this kind is known to arise in the quasiperiodic model at certain integer distances. For example, Ref. [37] observed that single-particle resonances are enhanced at orbitals separated by distances equal to the Fibonacci numbers $F_{n}$, when one picks $k=\phi$. Similarly, Ref. [76] reported suppressed interaction between LIOMs separated at the Fibonacci numbers for $k=\phi$, and the corresponding integer sequences that relate to the silver and bronze ratios. In our case, we find a more general condition that explains these anomalous system sizes, encompassing previous observations: they occur when the underlying potential is nearly commensurate with the lattice. For open boundary conditions, this corresponds to when $k(L+1)\approx n$ for some $n\in\mathbb{Z}$, so $k\approx\frac{n}{L+1}$ for some integer $n$. For example, in the calculation presented in Fig. 2(c), which was done for $k=\sqrt{2}$, there is a notable spike at $L=40$, and $\sqrt{2}\times 41\approx 58$. This observation helps explain why corresponding integer sequences arise for a given irrational number $k$, an example being the Fibonacci numbers when one chooses $k=\phi$, the golden ratio. Here, we note that $41$ is a member of the integer sequence $\frac{1}{2}Q_{n}$, where $Q_{n}$ are the Pell-Lucas numbers. Just as successive ratios of the Fibonacci numbers form the best rational approximations for $\phi$ ($\lim_{n\rightarrow\infty}F_{n+1}/F_{n}=\phi$), similarly, $\lim_{n\rightarrow\infty}\frac{1}{2}Q_{n}/P_{n}=\sqrt{2}$, where $P_{n}$ here refers to the Pell numbers. Thus, if $L+1=\frac{1}{2}Q_{n}$ for some $n$, then $(L+1)\times\sqrt{2}\approx m\in\mathbb{Z}$, leading to the potential being nearly commensurate with the lattice. We confirm that for $k=\phi$, these spikes occur at system sizes corresponding to the Fibonacci numbers and other integers $m$ that happen to satisfy $m\times k\approx n$. This underlying commensurability is likely the culprit that leads to increased resonances and other strange behavior in the spectrum, leading to a spike in the quantities calculated here. This would be expected as a periodic system exhibits vastly different physics from an incommensurate one.
@@ III.4.1 LIOM Corrections: Anderson Model
Figure 3: Analysis of the Frobenius norm of the first-order corrections to the noninteracting LIOMs, $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$, in the weakly interacting Anderson model. Each plot displays all orbitals $\alpha=1,\dots,L$ across the 500 disorder realizations used to generate the data shown in this figure. (a)-(d): Distributions of $\ln(\|\tilde{n}_{\alpha}^{(1)}\|_{F})$ at $W=0.6,1.6,2.6,3.1$ and for $\alpha=1,\dots,L$. Within each panel, each colored curve represents a different system size. We show calculations between $L=20$ and $L=100$ in steps of $10$. The inset in (d) shows the same data as in the main (d) panel, but with the vertical scale shown on log-scale. The red line in the inset represents a polynomial $\|\tilde{n}_{\alpha}^{(1)}\|_{F}^{-2}$ fit to the tail of the probability distribution of $\|\tilde{n}_{\alpha}^{(1)}\|_{F}$— see the text surrounding Eq. (24) for details on the form plotted. The convergence of the distributions for large enough $W$ and $L$ indicates the presence of an extensive number of quasiconservation laws after interactions are added, even in the thermodynamic limit. (e): Median $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ as a function of $L$ on a log-log scale. Each colored curve represents a different disorder strength $W$, from $W=0.1$ to $W=3.1$ in steps of $0.5$. One observes that the curves appear to saturate for large $L$, indicating that at least half of the LIOMs are conserved to second order when interactions are added. See Sec. III.4.1 for a full discussion of the results.
@@
We present numerical calculations of the Frobenius norm of the LIOM corrections, $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$, in Fig. 3, for the weakly–interacting Anderson model. For each $L$, we sample $500$ disorder realizations and record $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ for every orbital $\alpha=1,2,...,L$ in each disorder realization, giving a total of $500\times L$ datapoints in each distribution.
@@
The data in Fig. 3(a)–(d) suggest that the LIOM corrections converge to a fixed distribution in the thermodynamic limit. At low disorder, finite-size effects play a role, and the eigenstates appear extended. In these cases, the distributions initially shift rightward, but appear to converge at large $L$. At high disorder, the wavefunctions are localized and the distributions are converged, even at low $L$. Additionally, Fig. 3(e) shows how the median of these distributions grows with $L$. Here, we find that at low disorder, the median of $\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ grows as a polynomial with $L$. At higher disorders, we see similar growth of $\mathrm{med}\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ for low $L$, but the curves appear to have practically saturated with increasing system size. For example, for $W=3.1$, the median of the LIOM correction norms only increased from $1.37$ to $2.11$ between $L=20$ and $L=100$. This, combined with the fact that the growth of the curves of $\mathrm{med}\|\tilde{n}^{(1)}_{\alpha}\|_{F}$ is clearly slowing down significantly at larger $L$, suggests that at least half of the noninteracting LIOMs admit bounded corrections in the thermodynamic limit for high enough disorder. This is also consistent with the fact that the LIOM correction distributions have sufficiently converged at $L=100$ for strong enough disorder.
@@
Furthermore, in Fig. 4, we note that the $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ vs $L$ curves collapse when the x-axis is appropriately scaled by the participation ratio of the Anderson model, which scales approximately as $\mathrm{PR}(W)\propto W^{-1.78}$ in the range of $W$ considered here (see Appendix E for a more detailed explanation) and when the y-axis is scaled phenomenologically with $\sim\mathrm{PR}^{1.7}$. Here, the participation ratio is used as a proxy [^4] Some caution is needed in using the median participation ratio as a proxy for the localization length in this analysis. Firstly, both the participation ratio and the localization length depend on the energy of the noninteracting orbital, while our analysis does not consider the energy resolution of the LIOM corrections. Moreover, the power-law dependence of the localization length on the disorder strength is expected not to hold at large $W$, where a more refined ansatz is needed [77]. See Appendix E for another discussion. for the localization length $\xi$. The collapse of these curves suggests that at low disorders, the growth of $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is a finite-size effect. If the system size $L$ exceeds the localization length, for any disorder strength $\mathrm{med}\|\tilde{n}_{\alpha}^{(1)}\|_{F}$ is expected to saturate to a finite value. Put together, these results suggest that even in the thermodynamic limit, a fraction of noninteracting LIOMs have finite and small corrections in the presence of interactions. These LIOMs will survive for long times, supporting that thermalization can only happen on anomalously slow timescales in these systems: The rate of change of these approximate LIOMs can be rigorously bound by $\lambda^{2}\|[\hat{V},\hat{n}_{\alpha}^{(1)}]\|$, and a less rigorous, FGR-based estimate suggests a thermalization timescale of at least $O(\lambda^{-4})$ (see Sec. VII and Appendix F in Ref. [54]).
@@
Just as in the distributions of $\|\hat{X}\|_{F}$, the long $\|\hat{X}\|_{F}^{-2}$ power-law tails appear here for the same reason as they did in the AGP distributions [see the inset in Fig. 3(d)]. This shows that the perturbation theory breaks down for a fraction of the LIOMs, which receive large corrections due to resonances.